OpenStudy (armyrangers):

A circle has its center at (-1, 2) and a radius of 3 units. What is the equation of the circle? (x - 1)2 + (y + 2)2 = 3 (x + 1)2 + (y - 2)2 = 3 (x + 1)2 + (y + 2)2 = 9 (x + 1)2 + (y - 2)2 = 9

OpenStudy (armyrangers):

@BritBratt13

OpenStudy (anonymous):

(x-h)^2 + (y-k) ^2 = r^2 Standard equation of a circle

OpenStudy (mathmale):

When the center of the circle is at (0,0), and the radius is r, the standard equation is\[x^2+y^2=r^2\]and when the center is at (h,k), the equation is\[(x-h)^2+(y-k)^2 = r^2\]

OpenStudy (mathmale):

Please apply the appropriate formula to answering your posted question.

OpenStudy (armyrangers):

I don't know what I am doing

OpenStudy (mathmale):

You're given the center of the circle. What is it? See the original problem statement.

OpenStudy (armyrangers):

(-1,2

OpenStudy (mathmale):

OK: That point is your (h,k). the x-coordinate of the center is x=-1, and the x-coordinate is ... ???

OpenStudy (armyrangers):

-1

OpenStudy (mathmale):

That's right: h=-1. Next, k=?

OpenStudy (armyrangers):

K=2

OpenStudy (mathmale):

Good. Now, taking the standard form of the equation of a circle centered at (h,k), and substituting (-1) for h and substituting (2) for k, what do you get?\[(x-h)^2+(y-k)^2 = r^2\]

OpenStudy (armyrangers):

(X-(-1))^2+(y-2)^2=3^2

OpenStudy (mathmale):

Yes. Please simplify that. x-(-1)=? and so on.

OpenStudy (armyrangers):

(X+1)^2+(y-2)=9

OpenStudy (mathmale):

Looks great. Any questions about the procedures we've used here?

OpenStudy (armyrangers):

No thanks for the help

OpenStudy (mathmale):

Thank you for the medal. Hope to work with you again soon.